OpenStudy (anonymous):

What is an equation in point-slope form for (18, 26) and (20, 32)? What is the domain and range of the equation? Please explain the clearest and shortest way possible. Thanks!

OpenStudy (saifoo.khan):

Point slope form = \[y - y_1 = m(x-x_1)\] FIrst of all find the slope.

OpenStudy (saifoo.khan):

Second, insert the values of any one point, like (18, 26) or (20,32)

OpenStudy (saifoo.khan):

And that's it

OpenStudy (anonymous):

Thanks, but what's the equation? :)

OpenStudy (saifoo.khan):

do u want the direct answer? :O

OpenStudy (anonymous):

Kinda with the explanation + domain and range. lol

OpenStudy (saifoo.khan):

Sorry, i have no idea about domain and range.

OpenStudy (saifoo.khan):

\[y-26 = 3 (x-18) \]

OpenStudy (saifoo.khan):

Slope = 3

OpenStudy (anonymous):

Thanks :)

OpenStudy (saifoo.khan):

np